Help Needed;
Perform row operations on the augmented matrix as far as necessary to determine whether the system is independent, dependent, or inconsistent.
-x - y - z = 10
x + y + z = -6
x + y - z = 2
A) Inconsistent
B) Dependent
C) Independent

Answer by Alan3354(69443)

Perform row operations on the augmented matrix as far as necessary to determine whether the system is independent, dependent, or inconsistent.
-x - y - z = 10
x + y + z = -6
x + y - z = 2
A) Inconsistent
B) Dependent
C) Independent
------------
they're inconsistent since -1 times the first gives x+y+z=-10, and the 2nd says it's =6.
I determined that by inspection, so it wasn't necessary to do any row operations.